OpenStudy (anonymous):

Can someone tell me if I'm right or wrong? What is the solution set of |x – 2| + 4 = 3? There's no solution to this.

OpenStudy (johnweldon1993):

You would be correct

OpenStudy (anonymous):

Why do you say there are no solutions?

OpenStudy (primeralph):

@mrbarry tried solving it?

OpenStudy (anonymous):

I was trying to encourage thought, but someone said yes before I hit enter so I looked a bit dumb. I know the answer in my head.

OpenStudy (anonymous):

Yes, I already tried solving it. My answers were... x-2+4=3 +2 +2 +2 x+6=5 -6 -6 x= -1 then x-2-4=-6 +2 +2 +2 x-2-1 +2 +2 x= 1 Wait... So, I made a mistake in my previous calculation I just realized it.. The answer is -1 or 1. Right or wrong?

OpenStudy (anonymous):

Absolute value is always _________. Remember, to start the problem get the abs value isolated.

OpenStudy (anonymous):

I honestly have no idea what I'm doing Mr.Barry.... I'm just going with the flow and hoping I have right answers. I just need to know which one of my answers was right... ;~; Cause, you can explain all you like, I have NO idea what I'm doing.

OpenStudy (raden):

|x – 2| + 4 = 3 for x = -1 ----> |-1 – 2| + 4 = 3 |– 3| + 4 = 3 3 + 4 = 3 7 = 3 LHS not equal RHS

OpenStudy (anonymous):

Sure you do. If you start by getting the abs value isolated you get abs value = -1. That isn't possible so you were right. I was warned not to give away answers. Sorry to drag things out for you.

OpenStudy (anonymous):

Oh, so there was not solution! Thank goodness, I was right. xD

OpenStudy (noelgreco):

You're working way too hard. Leave the absolute value expression alone. I defer to mrbarry. Too many quick answers around here

OpenStudy (jhannybean):

|x-2|+4=3 |x-2| =-1 x - 2 = -1 OR -( x - 2 )= -1 x =-1 + 2 OR -x + 2 = -1 x = 1 OR -x = -3 => x=3 TO CHECK IF THESE ARE SOLUTIONS: X = 1 OR x = 3 |1 - 2| + 4 = 3 OR |3 - 2| + 4 = 3 |-1| + 4 = 3 OR |1| + 4 = 3 1 + 4 = 3 1 + 4 = 3 5 does not = 3 5 does not = 3